# SoK: LLM-based Log Parsing

### v1

@@ 3.1. Search Strategy
Subsequently, this section describes our search strategy and the inclusion and exclusion criteria derived from the core concepts of Sec. 2 and early insights from LLM-based log parsing approaches. For the keyword search, we determined the following keywords:
@@
  • It only applies an existing approach without introducing any novelties to LLM-based log parsing.
@@
  • It is not accessible with our institutions’ licenses.
By applying these exclusion criteria on both abstract and full text, we omitted $147$ of the initial publications found: $1$ publication was omitted because it was not written in English. $8$ publications were omitted due to being completely unrelated to the topic. $110$ are related but excluded, as they do not directly cover LLM-supported log data parsing by the definitions of Sec. 2 or are about other log analytics tasks, but employ parsing-free methods or existing conventional parsers, such as Drain (He et al., 2017) or SPELL (Du and Li, 2016). $21$ were omitted because they were books, theses, dissertations, or technical reports. $2$ were excluded for not being available freely with any license from our institutions. $5$ were outdated versions of newer papers.
The final selection consists of $29$ papers. One publication is from the year 2023, one from 2025, while the remaining are all from 2024. Since there is a recent hype around the usage of LLMs for log related issues, we include pre-print papers to capture the newest research. There are $10$ preprint papers in the final selection ($\sim 34\%$), all from $2024$.
@@ 3.2. Feature Definition
This section describes each feature and its classes of Table 2 which we report from the analyzed methods. We came up with this specific set of features, since they map to our research questions and pop up significantly often in many of the analyzed works. Additionally, we think that these are the most relevant for anyone who wants to apply LLM-based log parsing. The features either describe general properties (GP), processing steps (PS), or concern the reproducibility (R) of the approaches. For reproducibility, we take inspiration from the work of Olszewski et al. (Olszewski et al., 2023) where they analyzed machine learning papers of Tier 1 security conferences for reproducibility. The researchers defined a set of questions to determine the reproducibility of research findings.
In Table 2 features are reported with “✓” (true), an empty cell (false) or a descriptive name or abbreviation. If a paper provides an unclear answer to the feature in question, we write “?”. For instance, one papers provides a link to its code repository but the repository is empty, thus we write “?” for the corresponding feature. If a paper does not include the feature, or if the feature is not applicable, we leave it blank.
@@
In log parsing a labeled log is one for which a template, acting as the ground truth for that log, is available. We differentiate approaches based on the requirement for labels:
@@
Many methods are described as online approaches, but their interpretations vary. Some consider online processing to be incremental processing or streaming, while others apply batch-wise processing yet still label their method as online. Additionally, some works do not specify whether their approach is online or offline. Our initial incentive was to classify these methods accordingly, but this is often not feasible due to a lack of context or sufficient explanations. As a consequence, we devised the following categories that describe how many logs are processed at once:
  • Stream (S): The logs are processed one-by-one.
  • Batch (B): Multiple logs are processed at once, whereby it is possible to apply local (within each batch) optimizations. A batch is significantly smaller than the whole dataset.
@@
We report four different types of learning or prompt engineering techniques. The type of learning can strongly influence how the prompt is designed which is why we report this in a single feature:
@@
  • Fine-tuning (FT): In addition to ICL and CoT, which are solely modifications of the prompt, fine-tuning (Gao et al., 2021) modifies the parameters of the LLM, but mostly affects only the last few layers of the neural network, which are usually the most decisive ones for the outcome.
  • Pre-training (PT): Pre-training refers to the initial phase of training the model on a large corpus of text data to learn the structure and patterns of language.
@@
Some approaches increase their efficiency by storing parsed logs’ templates in some kind of data structure and only call the LLM if a subsequent log does not fit any of the existing templates. These data structures are mostly tree-like structures or similarity-based clusters.
@@
  • LogHub-2.0 (L2): LogHub-2.0 (Jiang et al., 2024b) is based on Loghub and features $14$ large-scale datasets from various domains with numbers of log in the order of $10^{4}$ to $10^{7}$.
@@
  • Edit Distance (ED): ED is the minimum number of operations, such as insertions, deletions, or substitutions, needed to convert one string into the other (Nedelkoski et al., 2021).
@@ 4. Literature Review Results
Astekin et al. (Astekin et al., 2024a) un, sup S ZS, S dir CL GA, ED, other GPT, Claude, Llama
@@
Ji et al. (Ji et al., 2024) (SuperLog) sup ? PT, ZS? dir L GA, other Llama, GPT, Qwen, OWL
@@
Pang et al. (Pang et al., 2024) (ONLA-LLM) un, sup S FT; ZS dir CU other Llama
@@
Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) un B S dir CA other OpenChat, Mistral, WizardLM
@@
Xiao et al. (Xiao et al., 2024) (LogBatcher) un B ZS S dir L2, CL GA, PA, ED GPT
@@ 4.1. Process Pipeline
After reviewing the paper selection, we were able to develop a process pipeline that summarizes all the approaches in a single flow chart, given in Fig. 2. The dashed arrows and boxes represent optional components while the continuous ones represent essential components. The user and the LLM represent key actors that can act on other components. The arrows pointing away from the user indicate steps that can require manual effort. The arrow from user to the logs represent the manual labeling effort for supervised parsers while the arrow pointing to the preprocessing component represents manual configuration. We provide detailed explanations about the components in the following sections.
@@ 4.2. Supervision (GP-1)
From the reviewed papers $6$ are fully supervised and $9$ approaches describe a supervised and an unsupervised setting. Supervised approaches require labeled logs for ICL or fine-tuning, but the amount of templates varies from a few guiding seed examples (Zhong et al., 2024) to significant proportions of a dataset’s templates (Jiang et al., 2024a; Xu et al., 2024). The works that fine-tune LLMs (Ma et al., 2024a; Zhong et al., 2024; Pang et al., 2024; Karanjai et al., 2024; Mehrabi et al., 2024) all require logs and their templates for the process which is why we classify these approaches as supervised.
@@ 4.3. Processing Mode (GP-2)
Log parsers can be categorized based on their processing mode, which determines whether they operate online or offline. The $18$ parsers classified under the “stream” category process logs in real-time as they arrive, making them suitable for online processing. These parsers handle data incrementally, allowing for immediate analysis and response which may be vital for downstream tasks using live monitoring. In contrast, “batch” parsers (Liu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024; Vaarandi and Bahsi, 2024) process logs in fixed-sized chunks, meaning they can function in both online and offline settings, depending on the frequency of batch execution. When processing in batches, it is possible to optimize the process for this batch (Xiao et al., 2024). Lastly, parsers under the category “total” (Zhang et al., 2025; Yu et al., 2024; Huang et al., 2024) analyze an entire log dataset at once, which is inherently an offline approach, as it requires the complete dataset to be available before processing begins. For instance, LUNAR (Huang et al., 2024) and Lemur (Zhang et al., 2025) group the logs based on certain features (e.g. length or most frequent shared oken) into buckets. They then create clusters within that buckets that maximize the sample diversity. The concept of maximizing the sample diversity within certain similarity clusters is also used by LogBatcher (Xiao et al., 2024) but for data batches. Similarly, DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this idea for their sampling algorithm.
@@ 4.4.2. Chain of Thought
Lemur (Zhang et al., 2025) uses CoT (Wei et al., 2022) to revise similar templates that may belong to the same event in three dialogue rounds. These steps include revising the structure, the semantics and finally, a decision if the templates should be merged. The other works applying CoT (Sun et al., 2024b, a; Xu and Gau, 2024) use it for direct parsing but do not explain in detail how.
@@ 4.4.3. Fine-Tuning and Pre-Training
From our selection only Ji et al. (SuperLog) (Ji et al., 2024) pre-trained a model. Pre-training requires vast amounts of data. To this end, Ji et al. created NLPLog, a comprehensive dataset with a quarter million question-answer pairs on log analysis tasks. The researchers report superior performance not only in parsing but also anomaly detection, failure diagnosis, and log interpretation. They also conducted an evaluation on unseen logs, but they do not report any of the common parsing metrics from Sec. 3.2. This hinders the determination of the practical benefits in settings involving unseen logs. Future research could be conducted on this aspect.
@@ 4.5. Manual Configuration (PS-1)
The datasets from LogHub and its descendants (Zhu et al., 2019, 2023; Khan et al., 2022) have been used for the evaluation of log parsing for many years now and it is known from existing literature which regular expressions, log formats, or other parameters have to be used to achieve a certain baseline of performance. Some of the reviewed methods require multiple parameters that are tuned specifically for each dataset and claim to achieve superior performance on the evaluation metrics. These matters may lead to unrealistic expectations for real-life applications, where sometimes no previous knowledge of the versatile format of logs is given (Dai et al., 2023). For instance, Dai et al. have shown that the parsing results of (non-LLM-based) parsers like Drain (He et al., 2017), IPLoM (Makanju et al., 2009), LenMA (Shima, 2016), AEL (Jiang et al., 2008) and Logram (Dai et al., 2022) are strongly influenced by their parameter settings. Of the reviewed studies, $9$ require certain configuration parameters. All of these $9$ require the log format of the logs to extract the log content from the raw logs and (Ma et al., 2024b; Zhang et al., 2025) additionally require regular expressions for preprocessing.
The requirement for manual configuration complicates a parser’s generic application for users and industry. Compared to the above mentioned list of non-LLM-based parsers, the reviewed LLM-based parsers exhibit less (essential) configuration parameters. This indicates that involving the LLM in the parsing process makes parsers more generic, thus reducing the need of human effort.
@@ 4.6. RAG (PS-2)
Ten approaches describe a sophisticated retrieval process, one describes random retrieval and for six approaches it is not clear how the approaches apply RAG. RAG is used for demonstration retrieval but also for retrieval of multiple target logs where a set of similar logs is chosen. The latter is only relevant for approaches parsing in batches or the whole dataset at once (GP-2). LogBatcher (Xiao et al., 2024), OpenLogParser (Ma et al., 2024b), LUNAR (Huang et al., 2024) prompt batches of logs that are in the same similarity cluster but maximize the dissimilarity within that cluster to highlight the variabilities and commonalities in logs for the LLM. Demonstrations can either be only logs (Karanjai et al., 2024) or logs and their templates (Xu et al., 2024; Jiang et al., 2024a; Pei et al., 2024; Zhong et al., 2024; Astekin et al., 2024a; Xiao et al., 2024; Ma et al., 2024b; Wu et al., 2024). The templates are either available through being parsed before, yet with no guaranty of correctness, or they origin from candidate sets of labeled logs for supervised settings (Xu et al., 2024; Jiang et al., 2024a; Sun et al., 2024b; Zhong et al., 2024; Pei et al., 2024)(GP-1).
@@
Another noteworthy phenomenon observed in generative language models is recency bias which refers to the tendency of the model to give disproportionate weight or attention to information that appears closer to the end of the prompt (Zhao et al., 2021). Consistently, Xu et al. (DivLog) (Xu et al., 2024) observed a significant impact on parsing performance based on the order of the retrieved examples in the prompt, whereby an ascending order of examples by similarity achieved the best results. The ordering of demonstrations in ascending order is adopted by multiple works (Xu et al., 2024; Wu et al., 2024; Xiao et al., 2024; Jiang et al., 2024a).
@@ 4.7. Caching (PS-3)
As LLM calls are costly compared to traditional algorithms, many approaches therefore create certain structures like clusters (Sun et al., 2024b; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2025; Jiang et al., 2024a; Karanjai et al., 2024; Huang et al., 2024; Pei et al., 2024), (template) trees (Ma et al., 2024b; Jiang et al., 2024a; Wu et al., 2024; Karanjai et al., 2024; Zhong et al., 2024; Karanjai et al., 2024) or other structures containing logs and their already parsed templates. In case the target log (or logs) matches an existing template (from an already parsed log) or is at least very similar, the parsing algorithm skips the LLM call and returns the corresponding template retrieved from the structure for this log. This self-evolutionary process (as Pei et al. (Pei et al., 2024) call it) is visualized in Fig. 2 by the arrow pointing from the templates to the cache component. This approach can greatly increase efficiency, allowing the usage of parsing with LLMs in settings where fast and cost efficient parsing is crucial, with reported runtimes comparable to runtime efficient parsers, like Drain (He et al., 2017). A parsing cache can also enhance effectiveness by mitigating the instability of results that often occurs with generative language models (Jiang et al., 2024a). Astekin et al. (Astekin et al., 2024b) find that even with temperature 00 some models do not answer deterministically. This emphasizes the usefulness of caches even more. However, it is possible that the first retrieved and stored template for an event type is incorrect, even if it matches some of the incoming logs. Such an event can be counteracted by a template revision step (PS-5). All but one paper (Sun et al., 2023), which described an approach with some form of caching, also described a post-processing step for correcting templates.
Since the total number of templates is drastically less than the number of logs, LLM parsing becomes a scalable approach with caching (Zhong et al., 2024). For example, the LogHub-2.0 datasets (Jiang et al., 2024b) contain more than $50$ million logs but less than $3500$ templates. $14$ approaches apply some form of caching, yet some do not explicitly name it “caching”. The structures used for caching can be efficiently and effectively used for sampling similar logs or relevant parsing demonstrations for ICL with RAG. Caching is not bound to the usage with LLMs and can be built into any other parsing approach.

### v2

@@ 4. Survey Method
This section describes the search strategy and the taxonomy of LLM-based log parsing.
@@ 4.1. Search Strategy
This section describes our search strategy and the inclusion and exclusion criteria derived from the core concepts of Sec. 2 and early insights from LLM-based log parsing approaches. For the keyword search, we determined the following keywords:
@@
  • It only applies a non-LLM-based approach without introducing novelties to LLM-based log parsing.
@@
By applying these exclusion criteria on both abstract and full text, we omitted $147$ of the initial publications found: $1$ publication was omitted because it was not written in English. $5$ were outdated versions of newer papers. $21$ were omitted because they were books, theses, dissertations, or technical reports. $8$ publications were omitted due to being completely unrelated to the topic. $112$ are related but excluded, as they do not directly cover LLM-based log parsing by the definitions of Sec. 2 or focus other log analytics tasks, but employ parsing-free methods or existing conventional parsers, such as Drain (He et al., 2017) or SPELL (Du and Li, 2016).
The final selection consists of $29$ papers. One publication is from the year 2023, one from 2025, while the remaining are all from 2024. Since there is a recent hype around the usage of LLMs for log related issues, we include preprint papers to capture the newest research. There are $12$ preprint papers in the final selection ($\sim 41\%$).
@@ 4.2. Feature Definition
This section describes each feature and its classes of Table 2, which we report from the analyzed approaches. The features describe either general properties (GP), processing steps (PS), or concern the reproducibility (R) of the approaches. Features concerning the general properties of the approaches are typical classifications of machine learning approaches. The features we categorize as processing steps are taken from the work of Jiang et al. (Jiang et al., 2024a) and Pei at al. (Pei et al., 2024). For reproducibility, we take inspiration from the work of Olszewski et al. (Olszewski et al., 2023) where they analyzed the reproducibility of machine learning papers of Tier 1 security conferences and defined a set of questions to determine the reproducibility of research findings. During the review process, we found that these features occur with significant frequencies or with significant variations, validating our selection. As illustrated in Table 2, this is evident.
In Table 2 features are reported with “✓” (true), an empty cell (false) or a descriptive name or abbreviation. If a paper provides an unclear answer to the feature in question, we write “?”. For instance, one papers provides a link to its code repository but the repository is empty, thus we write “?” for the corresponding feature. If a paper does not include the feature, we leave it blank.
@@
In log parsing, a labeled log is one for which a template, acting as the ground truth for that log, is available. In machine learning approaches are therefore often classified based on the requirement for labels into supervised (Wu et al., 2024; Liu et al., 2022) or unsupervised (Xiao et al., 2024; He et al., 2017) approaches, thus we report:
@@
Many methods are described as online approaches (He et al., 2017; Liu et al., 2024), but their interpretations vary. Some consider online processing to be incremental processing or streaming, while others apply batch-wise processing, yet still label their method as online. Additionally, some works do not specify whether their approach is online or offline. Our initial incentive was to classify these methods accordingly, but this is often not feasible due to a lack of context or sufficient explanations. As a consequence, we devised the following categories that describe how many logs are processed at once:
  • Stream (S): The log lines are processed one-by-one.
  • Batch (B): Multiple log lines are processed at once, whereby it is possible to apply local (within each batch) optimizations. A batch is significantly smaller than the entire dataset.
@@
We identified four different types of learning or prompt engineering techniques. The type of learning can strongly influence how the prompt is designed which is why we report this in a single feature:
@@
  • Fine-tuning (FT): In addition to ICL and CoT, which are solely modifications of the prompt, fine-tuning (Gao et al., 2021) modifies the parameters of the LLM, but mostly affects only the last few layers of the neural network, which are the most decisive ones for the outcome.
  • Pretraining (PT): Pretraining refers to the initial phase of training the model on a large corpus of text data to learn the structure and patterns of language.
@@
Some approaches increase their efficiency by storing parsed logs’ templates in some kind of data structure and only call the LLM if a subsequent log does not fit any of the existing templates. These data structures can be tree-like structures, similarity-based clusters, or simply lists.
@@
  • LogHub-2.0 (L2): LogHub-2.0 (Jiang et al., 2024b) is based on Loghub and features $14$ large-scale datasets from various domains with numbers of logs in the order of $10^{4}$ to $10^{7}$.
@@
  • Edit Distance (ED): ED is the minimum number of operations, such as insertions, deletions, or substitutions, needed to convert one string into the other (Nedelkoski et al., 2021). In log parsing, it is used to calculate the minimum number of operations required to convert the parsed template into the ground-truth template (Xiao et al., 2024).
@@
GA and PA are sensitive to the total number of log messages, which may be problematic since most log datasets contain a large number of logs but a much smaller number of unique templates (Jiang et al., 2024b). Therefore, Jiang et al. (Jiang et al., 2024b) proposed to use F-metrics since they are insensitive to class imbalance.
@@ 5. Literature Review Results
Astekin et al. (Astekin et al., 2024a) un, sup S ZS, S dir CL PA, ED, other GPT, Claude, Llama
@@
Ji et al. (Ji et al., 2024) (SuperLog) sup ? PT, ZS? dir L GA, FTA Llama, GPT, Qwen, OWL
@@
Pang et al. (Pang et al., 2024) (ONLA-LLM) un, sup S FT; ZS dir CU PA, PTA, RTA Llama
@@
Vaarandi et al. (Vaarandi and Bahsi, 2024) (LLM-TD) un B S dir CA FTA, other OpenChat, Mistral, WizardLM
@@
Xiao et al. (Xiao et al., 2024) (LogBatcher) un B ZS S dir L2, CL GA, PA, other GPT
@@ 5.1. Process Pipeline
After reviewing the paper selection, we were able to develop a process pipeline that summarizes all the approaches in a single flow chart, given in Fig. 2. The dashed arrows and boxes represent optional actions or components while the continuous ones represent essential actions components. Optional components can be skipped. The user and the LLM represent key actors that can act on other components. The arrows pointing away from the user indicate steps that can require manual effort. The arrow from user to the logs represent the manual labeling effort for supervised parsers while the arrow pointing to the preprocessing component represents manual configuration (which can also affect other components). We provide detailed explanations about the components in the following sections.
@@ 5.2. Supervision (GP-1)
From the reviewed papers, $6$ are fully supervised and $9$ approaches describe a supervised and unsupervised setting. Supervised ones require labeled logs for ICL or fine-tuning, but the amount of templates varies from a few guiding seed examples (Zhong et al., 2024) to significant proportions of a dataset’s templates (Jiang et al., 2024a; Xu et al., 2024). The works that fine-tune LLMs (Ma et al., 2024a; Zhong et al., 2024; Pang et al., 2024; Karanjai et al., 2024; Mehrabi et al., 2024) all require logs and their templates for the process which is why we classify these approaches as supervised.
@@ 5.3. Processing Mode (GP-2)
The $18$ parsers classified under the “stream” category process logs in real-time as they arrive, making them suitable for online processing. On the other hand, “batch” parsers (Liu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024; Vaarandi and Bahsi, 2024) process logs in fixed-sized chunks, meaning they can function in both online and offline settings, depending on the frequency of batch execution. When processing in batches, it is possible to optimize the process for this batch (Xiao et al., 2024). Lastly, parsers under the category “total” (Zhang et al., 2025; Yu et al., 2024; Huang et al., 2024) analyze an entire log dataset at once, which is inherently an offline approach, as it requires the complete dataset to be available before processing begins. For instance, LUNAR (Huang et al., 2024) and Lemur (Zhang et al., 2025) group the logs based on certain features (e.g., length or most frequent shared token) into buckets. They then create clusters within that buckets that maximize the sample diversity. The concept of maximizing the sample diversity within certain similarity clusters is also used by LogBatcher (Xiao et al., 2024) but for data batches. Similarly, DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this idea for their sampling algorithm.
@@ 5.4.2. Chain of Thought
Lemur (Zhang et al., 2025) uses CoT (Wei et al., 2022) to revise similar templates, that may belong to the same event, in three dialogue rounds. These steps include revising the structure, the semantics and finally, a decision if the templates should be merged. The other works applying CoT (Sun et al., 2024b, a; Xu and Gau, 2024) use it for direct parsing but do not explain in detail how.
@@ 5.4.3. Fine-Tuning and Pre-Training
From our selection only Ji et al. (SuperLog) (Ji et al., 2024) pretrained a model. Pretraining requires vast amounts of data. To this end, Ji et al. created NLPLog, a comprehensive dataset with a quarter million question-answer pairs on log analysis tasks. The researchers report superior performance not only in parsing but also anomaly detection, failure diagnosis, and log interpretation. They also conducted an evaluation on unseen logs, but they do not report any of the common parsing metrics from Sec. 4.2. This hinders the determination of the practical benefits in settings involving unseen logs. Future research could be conducted on this aspect.
@@ 5.5. Manual Configuration (PS-1)
The datasets from LogHub and its descendants (Zhu et al., 2019, 2023; Khan et al., 2022) have been used for the evaluation of log parsing for many years now and it is known from existing literature which regular expressions, log formats, or other parameters have to be used to achieve a certain baseline of performance. Some of the reviewed methods require multiple parameters that are tuned specifically for each dataset and claim to achieve superior performance on the evaluation metrics. These matters may lead to unrealistic expectations for real-life applications, where sometimes no previous knowledge of the versatile format of logs is given (Dai et al., 2023). For instance, Dai et al. have shown that the parsing results of (non-LLM-based) parsers like Drain (He et al., 2017), IPLoM (Makanju et al., 2009), LenMA (Shima, 2016), AEL (Jiang et al., 2008) and Logram (Dai et al., 2022) are strongly influenced by their parameter settings. Of the reviewed studies, $12$ require certain configuration parameters. All of them require the log format of the logs to extract the log content from the raw logs and (Ma et al., 2024b; Zhang et al., 2025) additionally require regular expressions for preprocessing.
@@ 5.6. RAG (PS-2)
Ten approaches describe a sophisticated retrieval process, one describes random retrieval and for six approaches it is not clear how the approaches apply RAG. RAG is used for demonstration retrieval but also for retrieval of multiple target logs where a set of similar logs is chosen. The latter is only relevant for approaches parsing in batches or the whole dataset at once (GP-2). LogBatcher (Xiao et al., 2024), OpenLogParser (Ma et al., 2024b), and LUNAR (Huang et al., 2024) prompt batches of logs that are in the same similarity cluster but maximize the dissimilarity within that cluster to highlight the variabilities and commonalities in logs for the LLM. Demonstrations can either be only logs (Karanjai et al., 2024) or logs and their templates (Xu et al., 2024; Jiang et al., 2024a; Pei et al., 2024; Zhong et al., 2024; Astekin et al., 2024a; Xiao et al., 2024; Ma et al., 2024b; Wu et al., 2024). The templates are either available through being parsed before, yet with no guaranty of correctness, or they origin from candidate sets of labeled logs for supervised settings (Xu et al., 2024; Jiang et al., 2024a; Sun et al., 2024b; Zhong et al., 2024; Pei et al., 2024)(GP-1).
@@
Another noteworthy phenomenon observed in generative language models is recency bias, which refers to the tendency of the model to give disproportionate weight or attention to information that appears closer to the end of the prompt (Zhao et al., 2021). Consistently, Xu et al. (DivLog) (Xu et al., 2024) observed a significant impact on parsing performance based on the order of the retrieved examples in the prompt, whereby an ascending order of examples by similarity achieved the best results. The ordering of demonstrations in ascending order is adopted by multiple works (Xu et al., 2024; Wu et al., 2024; Xiao et al., 2024; Jiang et al., 2024a).
@@ 5.7. Caching (PS-3)
As LLM calls are costly compared to traditional algorithms, many approaches therefore create certain structures like clusters (Sun et al., 2024b; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2025; Jiang et al., 2024a; Karanjai et al., 2024; Huang et al., 2024; Pei et al., 2024), prefix trees (Ma et al., 2024b; Jiang et al., 2024a; Wu et al., 2024; Karanjai et al., 2024; Zhong et al., 2024; Karanjai et al., 2024) or other structures containing logs and their already parsed templates. In case the target log (or logs) matches an existing template (from an already parsed log) or is at least very similar, the parsing algorithm skips the LLM call and returns the corresponding template retrieved from the structure for this log. This self-evolutionary process (Pei et al., 2024) is visualized in Fig. 2 by the arrow pointing from the templates to the cache component. This approach can greatly increase efficiency, allowing the usage of parsing with LLMs in settings where fast and cost efficient parsing is crucial, with reported runtimes comparable to runtime efficient parsers, like Drain (He et al., 2017). A parsing cache can also enhance effectiveness by mitigating the instability of results that often occurs with generative language models (Jiang et al., 2024a). Astekin et al. (Astekin et al., 2024b) find that even with temperature 00 some models do not answer deterministically. This emphasizes the usefulness of caches even more. However, it is possible that the first retrieved and stored template for an event type is incorrect, even if it matches some of the incoming logs. Such an event can be counteracted by a template revision step (PS-5). All but one paper (Sun et al., 2023), which described an approach with some form of caching, also described a post-processing step for correcting templates.
Since the total number of templates is drastically less than the number of logs, LLM parsing becomes a scalable approach with caching (Zhong et al., 2024). For example, the LogHub-2.0 datasets (Jiang et al., 2024b) contain more than $50$ million logs but less than $3500$ templates. $14$ approaches apply some form of caching, yet some do not explicitly name it “caching” and describe it simply as template matching. The structures used for caching can be efficiently and effectively used for sampling similar logs or relevant parsing demonstrations for ICL with RAG. Caching is not bound to the usage with LLMs and can be built into any other parsing approach.